# Expressing and visualizing model uncertainty in Bayesian variable selection using Cartesian credible sets

### v2

@@ Abstract
Modern regression applications can involve hundreds or thousands of variables which motivates the use of variable selection methods. Bayesian variable selection defines a posterior distribution on the possible subsets of the variables (which are usually termed models) to express uncertainty about which variables are strongly linked to the response. This can be used to provide Bayesian model averaged predictions or inference, and to understand the relative importance of different variables. However, there has been little work on meaningful representations of this uncertainty beyond first order summaries. We introduce Cartesian credible sets to address this gap. The elements of these sets are formed by concatenating sub-models defined on each block of a partition of the variables. Investigating these sub-models allow us to understand whether the models in the Cartesian credible set always/never/sometimes include a particular variable or group of variables and provide a useful summary of model uncertainty. We introduce methods to find these sets that emphasize ease of understanding. The potential of the method is illustrated on regression problems with both small and large numbers of variables.
@@ 1 Introduction
There are two main classes of Bayesian approach. Firstly, global-local shrinkage priors encourage coefficients which explain little variation to be shrunk very close to zero (see e.g. Bhadra et al., 2019, for a review). Post-processing methods can be used to perform variable selection often by finding a sparse summary of the posterior distribution (Hahn and Carvalho, 2015; Piironen et al., 2020). Alternatively, spike-and-slab priors place positive probability on some regression coefficients being exactly zero (Mitchell and Beauchamp, 1988; George and McCulloch, 1997). This leads to a posterior distribution on all possible models defined by subsets of the variables. This approach has good theoretical properties (Castillo et al., 2015) and predictive performance (Porwal and Raftery, 2022). Recent work has provided efficient algorithms for use in high-dimensional problems involving tens of thousands of regressors, including Markov chain Monte Carlo (MCMC) methods such as Importance Tempering (Zanella and Roberts, 2019), ASI (Griffin et al., 2021), LIT (Zhou et al., 2022) and PARNI (Liang et al., 2022), or stochastic search methods such as SVEN (Li et al., 2023). We will follow the spike-and-slab approach using MCMC but the methods could be extended to approximations of the posterior from a spike-and-slab prior found using different computational methods such as stochastic search.
@@
The paper is organized as follows. Section 2 briefly reviews Bayesian variable selection. Section 3 introduces a method for constructing credible model sets with a factorized posterior. Section 4 discusses approximating the posterior distribution with a distribution factorized according to a partition of the variables, and choosing the partition to balance the fidelity of the approximation with the complexity of the credible set using an “ease-of-understanding” criterion. Section 5 describes visualization of the Cartesian credible sets. Section 6 applies the methods to a simulated example. Section 7 shows how the method can be used with a GWAS of Systemic Lupus Erythematosus using a logistic regression model, and a linear regression model with two-way interactions. Section 8 discusses. Appendices include proofs, additional details of the Bayesian methods used in the examples, and further results. Code to implement the method on MCMC samples is available from https://jimegriffin.github.io/website/.
@@ 2 Bayesian variable selection
Variable selection methods find a subset of $p$ possible variables which explain the variation in a response through a regression model. We refer to each possible combination of included variables as a model, which is represented by the $p$-dimensional binary inclusion vector $\gamma$ where the $i$-th variable is only included in the model if $\gamma_{i}=1$. Let $\Gamma=\{0,1\}^{p}$ be the space of all possible models. Bayesian approaches assume that a prior is placed on $\Gamma$ as well as any parameters of the regression model. Tadesse and Vannucci (2021) provide a book length review and discuss recent research directions. A main focus of interest is the posterior distribution of the models. In this paper, we do not make any assumptions about the form of regression model or priors used to construct this posterior and we will write it generically as $p(\gamma\mid\mbox{Data})$ for a model $\gamma\in\Gamma$. This is a discrete distribution over $2^{p}$ possible models and so is hard to visualize or even write down unless $p$ is very small.
@@
As discussed in the introduction, there are several methods used to summarize the posterior distribution including: marginal PIPs, the Maximum a posterior (MAP) model and the median model (Barbieri and Berger, 2004; Barbieri et al., 2021). Barbieri and Berger (2004) show that the median model is optimal for prediction in orthogonal and nested design (when inclusion variables are independent a posteriori and the PIPs completely characterize the posterior distribution) and can be easily calculated from MCMC output. If there is correlation between variables then the inclusion variables will be a posteriori correlated and the PIPs do not summarize all information in the posterior. Barbieri et al. (2021) discuss properties of the median model under correlated designs. This is a particular problem if the inclusion variables have high correlation which corresponds to multi-collinearity in the design matrix of possible variables. To illustrate this effect, consider Example 5.2.2 in George and McCulloch (1997) where data is simulated from a linear regression model with $n=180$ observations and $p=15$ covariates. The only non-zero coefficients correspond to variables to 1, 3, 5, 7, 8, 11, 12, and 13. There is strong multicollinearity between the following blocks of variables: 1 and 2; 3 and 4; 5 and 6; 7, 8, 9, and 10; 11, 12 13, 14, and 15. Appendix B provides full details of the data generation and prior settings used in the analysis.
@@ 3.1 Credible sets
Credible sets extend credible intervals, which are commonly used to summarize posterior distributions, to posteriors on discrete spaces. Let $\mathcal{C}$ be the set of all possible subsets of $\Gamma$ then $S\subset\mathcal{C}$ is a $100\,\lambda\%$ credible set if $p(S\mid\mbox{Data})\geq\lambda$ (Wang et al., 2020). The $100\,\lambda\%$ highest posterior density credible interval is the shortest possible $100\,\lambda\%$ credible interval and, similarly, we can define the $100\,\lambda\%$ highest posterior probability (HPP) credible set, which is the smallest $100\,\lambda\%$ credible set. The $100\,\lambda\%$ HPP credible set can be found by ranking the models ranked list of the models in decreasing probability,
$p\left(\left.\gamma^{(1)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma^{(2)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma^{(3)}\right|\mbox{Data}\right)\geq\dots\geq p\left(\left.\gamma^{(2^{p})}\right|\mbox{Data}\right)$
and finding the smallest $K$ such that $\sum_{k=1}^{K}p\left(\left.\gamma^{(k)}\right|\mbox{Data}\right)\geq\lambda$ then $\left\{\gamma^{(1)},\gamma^{(2)},\dots,\gamma^{(K)}\right\}$ is the HPP credible set. This approach is possible if $p$ is relatively small (roughly less than 30) and all posterior probabilities can be calculated or approximated analytically. However, in many cases, $p$ will be larger and can be substantially larger (for example, genomic data sets can have tens of thousands of potential variables). In these cases, it’s often hard to accurately estimate model probabilities unless MCMC sample sizes are very large, as the level of model uncertainty makes it difficult to visit all plausible models, (i.e. models with posterior probability greater than $\epsilon$ for $10<\epsilon<$ 1 where, typically, $\epsilon$ is small) using stochastic search or MCMC. Additionally, searching the space of credible sets $\mathcal{C}$ is daunting as it contains $12^{2^{p}}-$ 1 possible sets and the size of the HPP credible set can make it hard to write down or consider.
@@
Table 1: George and McCulloch (1997) example: the models in the 50% HPP credible set shown with their posterior probability.. The vertical lines show the blocks of strongly multicollinear variables.
@@ 3.2 Finding Cartesian credible sets with factorized posteriors
     $q_{k^{\star}}=\frac{p\left(\left.\kappa_{k^{\star}}^{\left(r_{k^{\star}}\right)}\right|\mbox{Data}\right)/\sum_{r=r_{k^{\star}}}^{2^{l_{i}}}p\left(\left.\kappa_{k^{\star}}^{\left(r_{k^{\star}}\right)}\right|\mbox{Data}\right)}{2^{l_{i}}-r_{k^{\star}}}$.
@@ 4 Cartesian credible sets with general posteriors
If we can fully evaluate the posterior distribution then this can be calculated analytically. If there is an MCMC sample $\gamma^{(1)},\gamma^{(2)},\dots,\gamma^{(N)}\sim p(\gamma\mid\mbox{Data})$ then we can define $\kappa^{(i)}_{j}$ to be the values of $\gamma^{(i)}$ restricted to the block $P_{j}$. A Monte Carlo approximation to the KL divergence can be used
@@
We consider two algorithms with tuning parameters to find the partition which defines the factorized approximation. The first method uses a thresholded version of the correlation matrix and the second method applies an agglomerative algorithm to the KL divergence. In the thresholding method, we define a parameter $10<\eta<$ 1 and the matrix $\alpha(\eta)$ by $\alpha_{i,j}(\eta)=\mbox{I}(\|D_{i,j}|>\eta)$ for $i,j=1,\dots,p$ where $D$ is the posterior correlation matrix of $\gamma$. It is straightforward to show that $\alpha(\eta)$ is a well-defined adjacency matrix. The partition, as a function of $\eta$, is defined as the connected components of the graph defined by the adjacency matrix $\alpha(\eta)$. Clearly, the block sizes become larger as $\eta$ decreases. In practice, we find the partition for a grid of values of $\eta$, $01>\eta_{1}>\eta_{2}>\dots,\eta_{J}>$ 0 and find the corresponding partitions $P^{(1)},\dots,P^{(J)}$ and $100\lambda\%$ Cartesian credible set $S^{(1)},S^{(2)},\dots,S^{(J)}$ using Algorithm 1. In the agglomerative method, we define $\mbox{KL}(P)$ for a partition $P$ to be the Kullback-Leibler divergence between the posterior distribution and the approximation $q$ calculated using the optimal values of the parameters for $P$. A sequence of partitions can be calculated by merging two blocks of the partition at the previous iteration (in a similar way to agglomerative clustering). The two blocks to be merged are chosen to be ones that lead to the largest change in $\mbox{KL}(P)$. We define $\eta_{i}$ to be the change in $\mbox{KL}(P)$ at the $i$-th iteration of the algorithm. This greedy approach will lead to the largest improvement in the approximation at each stage and so defines a sequence of partitions $P_{1},\dots,P_{L}$ which are increasingly close approximation of the posterior distribution indexed by $\eta_{1},\dots,\eta_{L}$. The algorithm is summarized in Algorithm 2.
@@
5 Visualizing the $100\lambda\%$ Cartesian credible set
The $100\lambda\%$ Cartesian credible set gives us a convenient way to visualize the posterior distribution of $\gamma$. Firstly, for each block in the partition, we can list the sub-models in the corresponding block credible interval. If a block only includes the null sub-model, ${\bf 0}_{i}=\underbrace{(0,\dots,0)}_{l_{i}\mbox{ times}}$, it’s excluded from the graph. To provide further information about the posterior distribution, we colour the sub-models according to their marginal PIP. We also colour the partition according the posterior probability that at least one variable in the block is included in the model which is $p(S_{i}-{\bf 0}_{i}\mid\mbox{Data})$. This colouring is useful for understanding whether the block provides evidence for the inclusion of one of these variables and allows us to address dilution effects (George, 2010), the splitting of posterior probability between models including one of a set of strongly correlated variables.
6 George and McCulloch (1997) Example
We find 50% Cartesian credible sets with simulated data generated using the set-up in Example 5.2.2 of George and McCulloch (1997) by applying the algorithms developed in Section 4.
Thresholding Method
Agglomerative Method
@@ 6 George and McCulloch (1997) Example
Figure 2: George and McCulloch (1997) Example with $M=2$. Black lines indicate the blocks of the partition.
The sets are shown in Figure 2. Both methods find partitions with similar blocks to those identified as having strong multicollinearity in Section 2. Both methods split $\{7,8,9,10\}$ into two blocks $\{7,8,10\}$ and $\{9\}$. This has no effect on the final result as variables 7 and 8 are excluded from the model and variables 9 and 10 are included. The thresholding method also splits $\{11,12,13,14,15\}$ into two blocks $\{11,12,14,15\}$ and $\{13\}$. This leads to a slightly larger set (32 models) compared with the Agglomerative Method (24 models). This compares to the smallest 50% credible set which has 15 models. The difference is small compared with the total number of models (32768).
The Cartesian credible sets allows us to identify the interchangeability of variables and the colouring allows us to understand the overall importance of a block of variables and the contribution of each sub-model to the overall importance. For example, $S_{1}=\{(0,1),(1,0)\}$ and the overall posterior probability of $S_{1}$ is close to 0.9, the sub-models in $S_{1}$ are $(0,1)$ and $(1,0)$ which have corresponding posterior probabilities 0.38 and 0.52. This can be interpreted as evidence that one of variable 1 or variable 2 should be included in the model but there is no evidence for preferring either variables (although, the posterior probability slightly favours variable 1). We can draw similar conclusions about the pairs variables 3 and 4, and variables 5 and 6. The interpretation of the variables 11 to 15 depends slightly on the method for finding the partition. Using the thresholding method, there is evidence to include only one of the pairs variables 11 and 12, or variables 14 and 15. There is no evidence to include or exclude variable 13. Using the agglomerative method, there is evidence to include at least one of the variables 11 to 15. Again, variables 11 and 12, and variables 14 and 15 act as pairs. The thresholding method gives a slightly simpler interpretation but also a larger set. Overall, these results give a more succinct expression of model uncertainty than the 50% HPP credible set in Table 1 and gives more information than the PIPs or median model.

### v3

@@ Abstract
Modern regression applications can involve hundreds or thousands of variables which motivates the use of variable selection methods. Bayesian variable selection defines a posterior distribution on the possible subsets of the variables (which are usually termed models) to express uncertainty about which variables are strongly linked to the response. This can be used to provide Bayesian model averaged predictions or inference, and to understand the relative importance of different variables. However, there has been little work on meaningful representations of this uncertainty beyond first order summaries. We introduce Cartesian credible sets to address this gap. The elements of these sets are formed by concatenating sub-models defined on each block of a partition of the variables. Investigating these sub-models allow us to understand whether the models in the Cartesian credible set always/never/sometimes include a particular variable or group of variables and provide a useful summary of model uncertainty. We introduce a method to find these sets that emphasizes ease of understanding and can be easily computed from Markov chain Monte Carlo output. The potential of the method is illustrated on regression problems with both small and large numbers of variables.
@@ 1 Introduction
There are two main classes of Bayesian approach. Firstly, global-local shrinkage priors encourage coefficients which explain little variation to be shrunk very close to zero (see e.g. Bhadra et al., 2019, for a review). Post-processing methods can be used to perform variable selection often by finding a sparse summary of the posterior distribution (Hahn and Carvalho, 2015; Piironen et al., 2020). Alternatively, spike-and-slab priors place positive probability on some regression coefficients being exactly zero (Mitchell and Beauchamp, 1988; George and McCulloch, 1997). This leads to a posterior distribution on all possible models defined by subsets of the variables. This approach has good theoretical properties (Castillo et al., 2015) and predictive performance (Porwal and Raftery, 2022). Recent work has provided efficient algorithms for use in high-dimensional problems involving tens of thousands of regressors, including Markov chain Monte Carlo (MCMC) methods such as Importance Tempering (Zanella and Roberts, 2019), ASI (Griffin et al., 2021), LIT (Zhou et al., 2022) and PARNI (Liang et al., 2022), or stochastic search methods such as SVEN (Li et al., 2023). We will follow the spike-and-slab approach using MCMC but the methods could be extended to approximations of the posterior from a spike-and-slab prior found using different computational methods such as stochastic search. The inclusion of these priors in a Bayesian hierarchical model allows variable selection on latent variables at different levels of the hierarchy and makes these methods attractive to practitioners.
@@
The paper is organized as follows. Section 2 briefly reviews Bayesian variable selection. Section 3 introduces a method for constructing credible model sets with a factorized posterior. Section 4 discusses approximating the posterior distribution with a distribution factorized according to a partition of the variables, and choosing the partition to balance the fidelity of the approximation with the complexity of the credible set using an “ease-of-understanding” criterion. Section 5 describes visualization of the Cartesian credible sets. Section 6 applies the methods to simulated examples. Section 7 shows how the method can be used in different regression problems including a GWAS of Systemic Lupus Erythematosus using a logistic regression model. Section 8 discusses. Appendices include proofs, additional details of the Bayesian methods used in the examples, and further results. Code to implement the method on MCMC samples is available from https://jimegriffin.github.io/website/.
@@ 2 Bayesian variable selection
Variable selection methods find a subset of $p$ possible variables which explain the variation in a response through a regression model. We refer to variable in the subset as an included variable and each possible combination of included variables as a model. Each model can be represented by the $p$-dimensional binary inclusion vector $\gamma$ where the $i$-th variable is only included in the model if $\gamma_{i}=1$. Let $\Gamma=\{0,1\}^{p}$ be the space of all possible models. Bayesian approaches assume that a prior is placed on $\Gamma$ as well as any parameters of the regression model. Tadesse and Vannucci (2021) provide a book length review and discuss recent research directions. A main focus of interest is the posterior distribution of the models. In this paper, we do not make any assumptions about the form of regression model or priors used to construct this posterior and we will write it generically as $p(\gamma\mid\mbox{Data})$ for a model $\gamma\in\Gamma$. This is a discrete distribution over $2^{p}$ possible models and so is hard to visualize or even write down unless $p$ is very small.
@@
As discussed in the introduction, there are several methods used to summarize the posterior distribution including: marginal PIPs, the Maximum a posterior (MAP) model and the median model (Barbieri and Berger, 2004; Barbieri et al., 2021). Barbieri and Berger (2004) show that the median model is optimal for prediction in orthogonal and nested design (when inclusion variables are independent a posteriori and the PIPs completely characterize the posterior distribution) and can be easily calculated from MCMC output. If there is correlation between variables then the inclusion variables will be a posteriori correlated and the PIPs do not summarize all information in the posterior. Barbieri et al. (2021) discuss properties of the median model under correlated designs. This is a particular problem if the inclusion variables have high correlation which corresponds to multi-collinearity in the design matrix of possible variables. To illustrate this effect, consider Example 5.2.2 in George and McCulloch (1997) where data is simulated from a linear regression model with $n=180$ observations and $p=15$ covariates. The only non-zero coefficients correspond to variables to 1, 3, 5, 7, 8, 11, 12, and 13. There is strong multicollinearity between the following blocks of variables: 1 and 2; 3 and 4; 5 and 6; 7, 8, 9, and 10; 11, 12 13, 14, and 15. Appendix C provides full details of the data generation.
@@ 3.1 Credible sets
Credible sets extend credible intervals, which are commonly used to summarize posterior distributions, to posteriors on discrete spaces. Let $\mathcal{C}$ be the set of all possible subsets of $\Gamma$ then $S\subset\mathcal{C}$ is a $100\,\lambda\%$ credible set if $p(S\mid\mbox{Data})\geq\lambda$ (Wang et al., 2020). The $100\,\lambda\%$ highest posterior density credible interval is the shortest possible $100\,\lambda\%$ credible interval and, similarly, we can define the $100\,\lambda\%$ highest posterior probability (HPP) credible set, which is the smallest $100\,\lambda\%$ credible set. The $100\,\lambda\%$ HPP credible set can be found by listing the $2^{p}$ possible models by decreasing probability,
$p\left(\left.\gamma_{(1)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma_{(2)}\right|\mbox{Data}\right)\geq p\left(\left.\gamma_{(3)}\right|\mbox{Data}\right)\geq\dots\geq p\left(\left.\gamma_{(2^{p})}\right|\mbox{Data}\right)$
and finding the smallest $K$ such that $\sum_{k=1}^{K}p\left(\left.\gamma_{(k)}\right|\mbox{Data}\right)\geq\lambda$ then $\left\{\gamma_{(1)},\gamma_{(2)},\dots,\gamma_{(K)}\right\}$ is the HPP credible set. This approach is possible if $p$ is relatively small (roughly less than 30) and all posterior probabilities can be calculated or approximated analytically. However, in many cases, $p$ will be larger and can be substantially larger (for example, genomic data sets can have tens of thousands of potential variables) or be part of a larger hierarchical model. In these cases, it’s often hard to accurately estimate model probabilities unless MCMC sample sizes are very large, as the level of model uncertainty makes it difficult to visit all plausible models, (i.e. models with posterior probability greater than $\epsilon$ for $10<\epsilon<$ 1 where, typically, $\epsilon$ is small) using stochastic search or MCMC. Additionally, searching the space of credible sets $\mathcal{C}$ is daunting as it contains $12^{2^{p}}-$ 1 possible sets and the size of the HPP credible set can make it hard to write down or consider.
@@
Table 1: George and McCulloch (1997) example: the models in the 50% HPP credible set shown with their posterior probability. The vertical lines show the blocks of strongly multicollinear variables.
@@ 3.2 Finding Cartesian credible sets with factorized posteriors
     $q_{k^{\star}}=\frac{p\left(\left.\kappa_{k^{\star}}^{\left(r_{k^{\star}}\right)}\right|\mbox{Data}\right)/\sum_{r=r_{k^{\star}}}^{2^{l_{i}}}p\left(\left.\kappa_{k^{\star}}^{\left(r\right)}\right|\mbox{Data}\right)}{2^{l_{i}}-r_{k^{\star}}}$.
@@ 4 Cartesian credible sets with general posteriors
If we can fully evaluate the posterior distribution then this can be calculated analytically. If there is an MCMC sample $\gamma^{(1)},\gamma^{(2)},\dots,\gamma^{(N)}\sim p(\gamma\mid\mbox{Data})$ then we can define $\kappa^{(i)}_{j}$ to be the values of $\gamma^{(i)}$ restricted to the block $P_{j}$. If the blocks are relatively small then these sub-model posterior probabilities can be well-approximated using MCMC samples. A Monte Carlo approximation to the KL divergence can be used
@@
We use an agglomerative algorithm based on the KL divergence between the posterior and the factorized approximation to choose the tuning parameters which define the partition. Let $\mbox{KL}(P)$ be a partition $P$ to be the Kullback-Leibler divergence between the posterior distribution and the approximation $q$ calculated using the optimal values of the parameters for $P$. A sequence of partitions can be calculated by merging two blocks of the partition at the previous iteration (in a similar way to agglomerative clustering). The two blocks to be merged are chosen to be ones that lead to the largest change in $\mbox{KL}(P)$. We define $\eta_{i}$ to be the change in $\mbox{KL}(P)$ at the $i$-th iteration of the algorithm. This greedy approach will lead to the largest improvement in the approximation at each stage and so defines a sequence of partitions $P_{1},\dots,P_{L}$ which are increasingly close approximation of the posterior distribution indexed by $\eta_{1},\dots,\eta_{L}$. The algorithm is summarized in Algorithm 2.
@@
5 Interpreting and visualizing the $100\lambda\%$ Cartesian credible set
The $100\lambda\%$ Cartesian credible sets and the approximation of the posterior distribution as a product of distribution over blocks of variables give us a convenient way to interpret the model uncertainty expressed by the posterior distribution of $\gamma$. The product structure implies that the inclusion of blocks of variables are independent and so each block can be interpreted as a separate effect in the regression model. We define the block PIP for the $i$-th block to be $p(S_{i}-{\bf 0}_{i}\mid\mbox{Data})$ (where -- represents set difference and ${\bf 0}_{i}=\underbrace{(0,\dots,0)}_{l_{i}\mbox{ times}}$), which is the posterior probability that at least one variable in the block is included in the model. We also define the modal sub-model for the $i$-th block to be the sub-model with highest marginal posterior probability in block $i$. It is natural to consider the importance of a block of variables by considering whether the block PIP is larger than a threshold (for example, 0.5). If all blocks are singletons, the $\gamma_{i}$’a are independent under the approximation and this leads to the median model if the threshold is chosen to be 0.5. After deciding which blocks contain important variables, we consider the marginal posterior distribution of the sub-models in the important blocks. If the block is a singleton, the block PIP is just the marginal PIP and so that variable can be considered important. If the block is not a singleton, the posterior distribution of sub-models in a block can range from being very concentrated on the modal sub-model to being spread over a range of sub-models. Correspondingly, our interpretation can range from considering only the modal sub-model to being unable to choose between all or several sub-models.
To illustrate this method of interpretation and how the sets can be visualised, we find the $50\%$ Cartesian credible set with simulated data generated using the set-up in Example 5.2.2 of George and McCulloch (1997), Recall that there are $p=15$ covariates. The only non-zero coefficients correspond to variables to 1, 3, 5, 7, 8, 11, 12, and 13. There is strong multicollinearity between the following blocks of variables: 1 and 2; 3 and 4; 5 and 6; 7, 8, 9, and 10; 11, 12 13, 14, and 15. Appendix B gives prior settings and the MCMC scheme.
PIPs
50% Cartesian credible set
@@ 5 Interpreting and visualizing the $100\lambda\%$ Cartesian credible set
Figure 2: George and McCulloch (1997) example: Top row: PIPs for each variable with a thicker line indicating variables included in the median model. Bottom row: $50\%$ Cartesian credible set with $M=2$. Black lines indicate the blocks of the partition.
A graphical representation of the $50\%$ Cartesian credible set is shown in the bottom row of Figure 2. In this graph, the top row shows the variables which are included in at least one sub-model in the set. Here, variables 7 and 8 are excluded. Black lines above the variable names show the block. In this case, the block are $\{1,2\}$, $\{3,4\}$, $\{5,6\}$, $\{10\}$, $\{9\}$, and $\{11,12,13,14,15\}$. These are similar to grouping of variables with strong multicollinearity identified in Section 2 with the exception that $\{7,8,9,10\}$ is split into two blocks $\{10\}$ and $\{9\}$ and $7$ and $8$ are completely removed. The boxes around the variable numbers are coloured by the block PIP. All blocks have high block PIPs. Therefore, we can conclude that at least one of the variables in each block is important in the regression.
The sub-models in each block credible set are listed below the variables in the block. For example the sub-models for the first block are $S_{1}=\{(0,1),(1,0)\}$. This suggests that either 1 or 2 are included (but not both). Each sub-model is coloured according to their marginal posterior probability ($p(\kappa_{i}\mid\mbox{Data})$ for a sub-model $\kappa_{i}$). These are 0.38 for $(0,1)$ and 0.52 for $(1,0)$. This can be interpreted as evidence that one of 1 or 2 should be included in the model but there is not strong evidence for preferring one variable over the other (although, the posterior probability slightly favours variable 1). We can draw similar conclusions about the pairs of variables: 3 and 4, and 5 and 6. For the block $\{11,12,13,14,15\}$, there are three plausible submodels. One of these is the true subset of important variables which includes 11, 12 and 13. However, the marginal posterior probability of this submodel is only 0.18 and the two other models have stronger or similar support (0.51 and 0.17). The 50% Cartesian credible set has 24 models compared to 15 models in the 50% HPP credible set and the 50% HPP credible set is a subset of the 50% Cartesian credible set. The difference is small compared with the total number of models (32768).
It’s useful to compare the 50% Cartesian credible set to the median model which includes 1, 3, 4, 5, 9, 10, 14 and 15. This model includes the correct number of variables (8) but includes 5 variables which are not actually important (4, 9, 10, 14, 15). The 50% Cartesian credible set provides a more nuanced summary. Firstly, the median model chooses both 3 and 4 when the posterior distribution shows clear support for one (but not both) being included. Secondly, the median model excludes the truly important variables 11, 12 and 13 and selects the (correlated) variables 14 and 15. The 50% Cartesian credible set shows that at least one of $\{11,12,13,14,15\}$ should be included and there are three plausible sub-models. One of these sub-models is the true combination of variables $(1,1,1,0,0)$. The median model is forced to choose a single model and ignores information about the relationship between variables. The Cartesian credible set approach allows us to visualize the relationship between variables and to concentrate on collections of plausible sub-models. This avoids the dangers of overinterpreting the median model (particularly for variables with PIPs close to 0.5) and can be used to find, for example, the interchangeability of variables and allows us to address dilution effects (George, 2010), the splitting of posterior probability between models including subsets of a set of strongly correlated variables.
6 Simulated example: the effects of sample size in block correlated example
To illustrate how the Cartesian credible set changes with sample size, we consider the following simulated example. Data are simulated from a linear regression with $p=15$ regressors and $n=640$ observations where
$y_{i}=X_{i}\beta+\epsilon_{i}$
with $\epsilon_{i}\sim\mbox{N}(0,\sigma^{2})$, $X_{i}\sim\mbox{N}(0,A(\rho))$ and $\beta=(1,-1,0,1,0,0,-1,0,0,0,1,0,0,0,0)$ (so that the 1st, 2nd, 4th, 7th and 11th regression coefficients are non-zero). The $p\times p$ correlation matrix of the regressors is block diagonal leading to correlation between regressors in the same block and has the form
$A(\rho)=\left(\begin{array}[]{ccccc}A_{1}(\rho)&0&\ldots&0\\ 0&A_{2}(\rho)&\ldots&0\\ \vdots&\vdots&\ddots&\\ 0&0&&A_{5}(\rho)\end{array}\right)$
where the $(j,k)$-th element in $A_{i}(\rho)$ is $\rho^{|j-k|}$and higher values of $\rho$ lead to stronger correlations between regressors in the same block. The choice of regression coefficients leads to one (and only one) non-zero regression coefficient in each block. We would expect that it is harder to distinguish the effects of variables in the same block as $\rho$ increases and that this effect decreases as $n$ increases. The prior settings and MCMC scheme used for inference are given in Appendix B.
We fit a linear regression model to the data using the first $n$ observations for $n=40$, $n=80$, $n=160$, $n=320$ and $n=640$.
$n$ PIP CCS
40 Refer to caption Refer to caption
80 Refer to caption Refer to caption
160 Refer to caption Refer to caption
320 Refer to caption Refer to caption
640 Refer to caption Refer to caption
Figure 3: Simulated example: PIPs and $50\%$ Cartesian credible sets (CCS) with $M=2$ for simulated datasets with $\rho=0.5$ and different sample sizes. The variables in the median model are indicated by thicker lines in the PIP plot.
The 50% Cartesian credible sets are presented in Figure 3 for $\rho=0.5$. As expected, the important variables are mostly clear for a sufficiently large value of $n$ ($n=80$ in this case). When $n=80$, 1, 2, 4, and 7 are included as singleton blocks which each have high PIPs. The other important variable, 11, is included in a block with regression coefficient 14 which expresses uncertainty about which variable is important. The median model includes both 11 and 14 but the graphical representation of the 50% Cartesian credible set shows that this is because the posterior probability that both variables are included is 0.53 and there is also support for only one of the variables being included. For $n=160$ and $n=320$, blocks of unimportant variables are identified but in total these blocks have block PIPs below 0.5 which suggests that these groups of variables are not important.
$n$ PIP CCS
40 Refer to caption Refer to caption
80 Refer to caption Refer to caption
160 Refer to caption Refer to caption
320 Refer to caption Refer to caption
640 Refer to caption Refer to caption
Figure 4: Example 2: PIPs and $50\%$ Cartesian credible sets (CCS) with $M=2$ for simulated datasets with $\rho=0.9$ and different sample sizes. The variables in the median model are indicated by thicker lines in the PIP plot.
The results with $\rho=0.9$ are shown in Figure 4. As expected, the larger level of correlation within blocks of variables makes the variable selection problem harder. The median model tends to choose too many variables. For example, 3, 9, and 15 are included in the median model when $n=80$ and 3, 8, 9, and 14 when $n=160$. The underselection disappears as $n$ increases and the five truly important variables are selected when $n=640$. This underselection is due to the correlation in the regressors leading to more than one variable in a block being selected. The 50% Cartesian credible set allows us to understand this. For example, when $n=80$, (2, 3), (7, 9) and (11, 14, 15) are indicated in blocks. This shows a truer picture of the data generating mechanism that one of 2 or 3, one of 7 and 9, and one 11, 14 and 15 should be included in the model. We see a similar effect when $n=160$, the blocks again show the relationships between variables 2 and 3, and 7, 8, 9. However, variable 14 is not included in a block with variable 11 since variable 11 has a high PIP. When $n=320$ and $n=640$, the posterior concentrates posterior probability on the true model (and very similar models) and so the ambiguity about the important variables disappears as again illustrated by the Cartesian credible sets.